OpenStudy (anonymous):

Joseph is building a cone using modeling clay. The cone has a radius of 7 cm and a height of 12.5 cm. Joseph adds additional clay, keeping the radius the same, until the cone reaches a height of 18 cm. How much clay did Joseph add? Use 3.14 to approximate pi and round the answer to two decimal places. ____cm3 @Directrix

Directrix (directrix):

Strategy: the taller cone volume minus the smaller cone volume

OpenStudy (anonymous):

so 18-7?

Directrix (directrix):

Taller: Vcone = 1/3 * 3.14 * 7 * 18 @TeenWolfGirl What is this: 1/3 * 3.14 * 7 * 18 =

OpenStudy (anonymous):

131.88

Directrix (directrix):

Shorter: Vcone = 1/3 * 3.14 * 7 * 18 @TeenWolfGirl What is this: 1/3 * 3.14 * 7 * 18 =

OpenStudy (anonymous):

131.88

Directrix (directrix):

131.88 Correct for Taller Volume

OpenStudy (anonymous):

either of you know the value of negative 29?

OpenStudy (anonymous):

how about you post the question

OpenStudy (anonymous):

this is my question because i need help

Directrix (directrix):

Whoa. I messed up on the shorter volume.

Directrix (directrix):

Shorter: Vcone = 1/3 * 3.14 * 7 * 12.5 Sorry. @TeenWolfGirl 1/3 * 3.14 * 7 * 12.5 = ?

OpenStudy (anonymous):

31.583 so now i subtract, correct?

Directrix (directrix):

31.583 is not the volume of the shorter cone. Or, I messed up.

OpenStudy (anonymous):

sorry no meant 91.583

OpenStudy (anonymous):

131.88-91.583=100.297

Directrix (directrix):

Okay, so do 131.88 - 91.58 for the final

Directrix (directrix):

I did not get what you did.

OpenStudy (anonymous):

oh then i would get 131.88-31.58=100.3

OpenStudy (anonymous):

so my answer is 100.3, correct?

Directrix (directrix):

It is this: 131.88 - 91.58 Not 131.88-31.58

OpenStudy (anonymous):

i meant 91.58

Directrix (directrix):

so my answer is 100.3, correct? NO

OpenStudy (anonymous):

40.3, correct?

OpenStudy (anonymous):

Thanks

Directrix (directrix):

Yes, 40.30 is what I got.

Directrix (directrix):

You are welcome.